# AAPO: Enhance the Reasoning Capabilities of LLMs with Advantage Momentum

### v1

@@ 4 Advantage-Augmented Policy Optimization
From the formulation of the augmented advantage $\hat{A}_{i,t}^{*}$, it is evident that as the policy is optimized, the reward $r_{\theta_{i}}$ of the responses sampled from the policy $\pi_{\theta}$ within a group increases. Consequently, the distance between these rewards $r_{\theta_{i}}$ and those $r_{ref_{i}}$ of the responses sampled from the reference model $\pi_{ref}$ will also widen, since the parameters of the reference model remain frozen throughout the training process of AAPO. In later steps of RL, the responses sampled from the policy tend to be of high quality, causing the relative advantages $\hat{A}_{i,t}^{GRPO}$ within the group to approach zero. If training continues using the original advantage estimation method in Equation (2), the resulting gradients will approach zero, leading to significantly reduced training efficiency. However, under the proposed method augmenting advantage with advantage momentum in Equation (5), even when the group relative advantage approaches zero, the rewards of the policy samples remain higher than those of the reference samples. This ensures a nonzero advantage $\hat{A}_{i,t}^{*}$ in AAPO as discussed in Section 5.2, thereby maintaining informative gradients for continued policy optimization.
@@
AAPO Training The training process of AAPO can be summarized as follows: 1) Sample two groups of responses, $O_{\theta}$ and $O_{ref}$, of equal size from both the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$, respectively. 2) For each response in $O_{\theta}$, compute its group relative advantage and the advantage momentum with the corresponding response in $O_{ref}$. The values, calculated by Equation (5), are then used to perform gradient updates according to Equation (6). It is important to note that throughout the training process, the parameters of the reference model $\pi_{ref}$ remain frozen and do not undergo gradient updates. The formal training procedure is described in Algorithm 1.
@@
8:        5) compute the augmented advantage $\hat{A}_{i,t}^{*}$ following Equation (5) for each sample in $O_{\theta}$
9:        6) compute loss for $O_{\theta}$ following Equation (4)
@@
11:     7) update $\pi_{\theta}$ following Equation (6)
@@
5 Analysis of AAPO
5.1 Deep Analysis of Group Relative Advantage Estimation
As shown in Equation (2), current advantage estimation methods in recent RL approaches [36, 48, 25] that eliminate the dependency on a value model predominantly adopt this form of computation. We now present a rigorous analysis into the underlying phenomena induced by this advantage estimation method: (1) Phenomenon 1: What are the implications when all responses within a group are similarly good (or all of high quality)? (2) Phenomenon 2: What are the implications when all responses within a group are similarly bad (or all of low quality)?
Phenomenon 1 Considering that all responses are similarly good, which implies that the reward for each response is nearly the same, this indicates $\forall i,j\in\{1,2,3,\cdots,G\}\land i\neq j,\text{ }r_{i}\approx r_{j}$, their respective advantage, as estimated according to Equation (2), can be expressed in the following form:
$\hat{A}_{i,t}^{GRPO}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}\to 0.$ (7)
As expressed in Equation (7), the advantage of each response approaches zero in this case. For illustrative purposes, we use the loss function of GRPO [48] as an example. As mentioned in DAPO [48], removing the KL divergence in GRPO could further improve the optimization. The KL divergence, summation and averaging operators in Equation (1) are omitted to facilitate clarity in this context. Under these conditions, the gradient formula of GRPO is formally given by:
$\nabla_{\theta}\mathcal{J}_{\mathrm{GRPO}}(\theta)=$ (8)
$\left\{\begin{array}[]{ll}\hat{A}_{i,t}^{GRPO}r_{i,t}(\theta)\nabla_{\theta}\log\pi_{\theta}\left(a_{t}\mid s_{t}\right),&f_{i,t}(\theta)<g_{i,t}(\theta).\\ \hat{A}_{i,t}^{GRPO}\left[\operatorname{clip}\left(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}}\right)\right]^{\prime}\nabla_{\theta}\log\pi_{\theta}\left(a_{t}\mid s_{t}\right),&g_{i,t}(\theta)\leq f_{i,t}(\theta).\end{array},\right.$
where
$f_{i,t}(\theta)=r_{i,t}(\theta)\hat{A}_{i,t}^{GRPO},\quad g_{i,t}(\theta)=\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon_{\mathrm{low}},1+\varepsilon_{\mathrm{high}}\right)\hat{A}_{i,t}^{GRPO}.$ (9)
As illustrated in the computation of Equation (8), when the advantage of each response approaches zero, the corresponding gradient also decreases to zero. This implies that the gradient update for the policy becomes negligible for this training step, resulting in very low training efficiency. Similar issues are observed in methods such as GRPO, DAPO, GPG [8], and Dr. GRPO [25]. Nonetheless, the observation that all responses are associated with similarly high reward does not unequivocally imply that the policy has been sufficiently optimized; it may alternatively reflect a high variance [32, 47] in the current policy, suggesting that the policy has converged to a sub-optimum, performing well only on a narrow category of questions. However, this phenomenon is frequently observed during the later steps of RL training if the advantage estimation method in GRPO is adopted.
Phenomenon 2 Similar to the scenario in Phenomenon 1 where all responses are similarly bad, when all responses are similarly bad, the reward assigned to each response tends to be similar. As a result, the relative advantage of each response approaches zero, leading to a zero gradient during policy updates, which is computed using Equation (8). Consequently, the efficiency of this training step becomes significantly low. Such a phenomenon is frequently observed when the policy is confronted with input samples that exhibit inherently high complexity or ambiguous representation.
Analysis Conclusion Based on our in-depth analysis of the Phenomenon 1 and the Phenomenon 2 where generated responses are similarly good and similarly bad within a group, we observed that the gradient tends to approach zero. This results in extremely low training efficiency during RL training.
Above phenomena can be generalized When the rewards of responses within any given group are identical or highly similar, regardless of whether the responses are all good or all bad, the gradient approaches zero, rendering the gradient update nearly ineffective in the training step. This issue becomes particularly pronounced in the later steps of RL training, where generated responses are consistently of high quality, and the corresponding advantage approaches zero. This indicates that the training efficiency progressively declines as RL training progresses. Motivated by this insight, we propose an advantage-augmented RL algorithm AAPO to address the aforementioned phenomena.
5.2 Understanding the effectiveness of AAPO
As discussed in Section 5.1, the previous advantage estimation method (as expressed in Equation (2)) in RL algorithms [36, 8, 25] can lead to the advantage value approaching zero, which in turn causes the magnitude of gradient updates to diminish accordingly. In this section, we provide a comprehensive analysis of why our proposed advantage-augmented method can effectively mitigate these issues.
Analysis 1 Consider a general situation, which naturally encompasses the Phenomenon 1 and the Phenomenon 2 in Section 5.1. In the later steps of RL training, once the policy has already acquired relatively easier-to-learn features, it often struggles to learn more complex ones. During this phase, when the policy $\pi_{\theta}$ samples a group of responses, the reward associated with each sample tends to be similar. Consequently, the relative advantage computed according to Equation (2) approaches zero. To address this issue, we propose AAPO, which estimates the advantage $\hat{A}_{i,t}^{*}$ with advantage momentum following Equation (5). Since the capability of the reference model $\pi_{ref}$ remains unchanged while the policy model $\pi_{\theta}$ improves progressively throughout AAPO training, the quality of responses generated by the policy model $\pi_{\theta}$ exceeds that of the reference model $\pi_{ref}$ over time. By measuring the distance between the rewards of two groups of responses $O_{\theta}=\{o_{\theta_{1}},o_{\theta_{2}},\cdots,o_{\theta_{G}}\}$ and $O_{ref}=\{o_{ref_{1}},o_{ref_{2}},\cdots,o_{ref_{G}}\}$ from $\pi_{ref}$ generated by the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$, respectively, we can calculate the augmented advantages of each response in $O_{\theta}$ via Equation (5). This prevents the advantages from approaching zero and ensures that the gradients used to update the policy remain informative and effective. This analysis can be generalized to any situation where the rewards of the responses in the group are similar, whether good or not.
Analysis 2  Group relative advantage estimation (Equation (2)) can be problematic in the presence of two types of asymmetric response distributions. In one case, where most responses from the policy $\pi_{\theta}$ are high quality except for a single outlier, the relatively worse response may receive a negative advantage and contribute an opposing gradient, increasing variance. Although this response may still be correct under multi-dimensional reward rules, such as both format and correctness, it is penalized due to its format. This misalignment between reward attribution and the true value of a response increases the risk of reward hacking[11, 28]. In the opposite case, where most responses are low quality and one is relatively better, the better one receives an excessively high advantage despite possibly low absolute quality, leading to biased updates and suboptimal convergence. Our method, which augments advantage estimation with advantage momentum and compares responses to the reference model $\pi_{\text{ref}}$, addresses both issues: it boosts underappreciated yet valuable responses and suppresses misleading ones that exhibit a high estimated advantage. As a result, AAPO reduces variance and the risk of reward hacking, and improves the optimization stability of the policy.
Analysis 3  When policy optimization with AAPO reaches a global optimum, which means $\forall i,j\in\{1,2,3,\cdots,G\}\land i\neq j,\text{ }r_{i}\approx r_{j}$, the objective in Equation (6), which consists of the CE loss and the augmented advantage, exhibits a specific and predictable behavior as detailed in our derivation below (for clarity and notational convenience, we omit the expectation operator):

### v2

@@ 4 Advantage-Augmented Policy Optimization
From the formulation of the augmented advantage $\hat{A}_{i,t}^{*}$, it is evident that as the policy is optimized, the reward $r_{\theta_{i}}$ of the responses sampled from the policy $\pi_{\theta}$ within a group increases. Consequently, the distance between these rewards $r_{\theta_{i}}$ and those $r_{ref_{i}}$ of the responses sampled from the reference model $\pi_{ref}$ will also widen, since the parameters of the reference model remain frozen throughout the training process of AAPO. In later steps of RL, the responses sampled from the policy tend to be of high quality, causing the relative advantages $\hat{A}_{i,t}^{\mathrm{GRPO}}$ within the group to approach zero. If training continues using the original advantage estimation method in equation 2, the resulting gradients will approach zero, leading to significantly reduced training efficiency. However, under the proposed method augmenting advantage with advantage momentum in equation 5, even when the group relative advantage approaches zero, the rewards of the policy samples remain higher than those of the reference samples. This ensures a nonzero advantage $\hat{A}_{i,t}^{*}$ in AAPO as discussed in Section 5.2, thereby maintaining informative gradients for continued policy optimization.
5 Analysis of AAPO
5.1 Deep Analysis of Group Relative Advantage Estimation
As shown in equation 2, current advantage estimation methods in recent RL approaches (Shao et al., 2024; Yu et al., 2025; Liu et al., 2025) that eliminate the dependency on a value model predominantly adopt this form of computation. We now present a rigorous analysis into the underlying phenomena induced by this advantage estimation method: (1) Phenomenon 1: What are the implications when all responses within a group are similarly good (or all of high quality)? (2) Phenomenon 2: What are the implications when all responses within a group are similarly bad (or all of low quality)?
Phenomenon 1 Considering that all responses are similarly good, which implies that the reward for each response is nearly the same, this indicates $\forall i,j\in\{1,2,3,\cdots,G\}\land i\neq j,\text{ }r_{i}\approx r_{j}$, their respective advantage, as estimated according to equation 2, can be expressed in the following form:
$\hat{A}_{i,t}^{\mathrm{GRPO}}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}\to 0.$ (7)
As expressed in equation 7, the advantage of each response approaches zero in this case. For illustrative purposes, we use the loss function of GRPO (Yu et al., 2025) as an example. As mentioned in DAPO (Yu et al., 2025), removing the KL divergence in GRPO could further improve the optimization. The KL divergence, summation and averaging operators in equation 1 are omitted to facilitate clarity in this context. Under these conditions, the gradient formula of GRPO is formally given by:
$\nabla_{\theta}\mathcal{J}_{\mathrm{GRPO}}(\theta)=A_{i,t}^{\mathrm{GRPO}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t}).$ (8)
As illustrated in the computation of equation 8, when the advantage of each response approaches zero, the corresponding gradient also decreases to zero. This implies that the gradient update for the policy becomes negligible for this training step, resulting in very low training efficiency. Similar issues are observed in methods such as GRPO, DAPO, GPG (Chu et al., 2025), and Dr. GRPO (Liu et al., 2025). Nonetheless, the observation that all responses are associated with similarly high reward does not unequivocally imply that the policy has been sufficiently optimized; it may alternatively reflect a high variance (Roelofs et al., 2019; Yu et al., 2022) in the current policy, suggesting that the policy has converged to a sub-optimum, performing well only on a narrow category of questions. However, this phenomenon is frequently observed during the later steps of RL training if the advantage estimation method in GRPO is adopted.
Phenomenon 2 Similar to the scenario in Phenomenon 1 where all responses are similarly bad, when all responses are similarly bad, the reward assigned to each response tends to be similar. As a result, the relative advantage of each response approaches zero, leading to a zero gradient during policy updates, which is computed using equation 8. Consequently, the efficiency of this training step becomes significantly low. Such a phenomenon is frequently observed when the policy is confronted with input samples that exhibit inherently high complexity or ambiguous representation.
Analysis Conclusion Based on our in-depth analysis of the two phenomena where generated responses are similarly good and similarly bad within a group, we observed that the gradient tends to approach zero. This results in extremely low training efficiency during RL training.
Above phenomena can be generalized When the rewards of responses within any given group are identical or highly similar, regardless of whether the responses are all good or all bad, the gradient approaches zero, rendering the gradient update nearly ineffective in training. This issue becomes particularly pronounced in the later steps of RL training, where generated responses are consistently of high quality, and the corresponding advantage approaches zero. This indicates that the training efficiency progressively declines as RL training progresses. Motivated by this insight, we propose an advantage-augmented RL algorithm AAPO to address the aforementioned phenomena.
5.2 Understanding the effectiveness of AAPO
As discussed in Section 5.1, the previous advantage estimation method (as expressed in equation 2) in RL algorithms (Shao et al., 2024; Chu et al., 2025; Liu et al., 2025) can lead to the advantage value approaching zero, which in turn causes the magnitude of gradient updates to diminish accordingly. In this section, we provide a comprehensive analysis of why our proposed advantage-augmented method can effectively mitigate these issues.
Analysis 1 Consider a general situation, which naturally encompasses the two phenomena in Section 5.1. In the later steps of RL training, once the policy has already acquired relatively easier-to-learn features, it often struggles to learn more complex ones. During this phase, when the policy $\pi_{\theta}$ samples a group of responses, the reward associated with each sample tends to be similar. Consequently, the relative advantage computed according to equation 2 approaches zero. To address this issue, we propose AAPO, which estimates the advantage $\hat{A}_{i,t}^{*}$ with advantage momentum following equation 5. Since the capability of the reference model $\pi_{ref}$ remains unchanged while the policy model $\pi_{\theta}$ improves progressively throughout AAPO training, the quality of responses generated by the policy model $\pi_{\theta}$ exceeds that of the reference model $\pi_{ref}$ over time. By measuring the distance between the rewards of two groups of responses $O_{\theta}=\{o_{\theta_{1}},o_{\theta_{2}},\cdots,o_{\theta_{G}}\}$ and $O_{ref}=\{o_{ref_{1}},o_{ref_{2}},\cdots,o_{ref_{G}}\}$ from $\pi_{ref}$ generated by the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$, respectively, we can calculate the augmented advantages of each response in $O_{\theta}$ via equation 5. This prevents the advantages from approaching zero and ensures that the gradients used to update the policy remain informative and effective. This analysis can be generalized to any situation where the rewards of the responses in the group are similar, whether good or not.
Analysis 2  Group relative advantage estimation (equation 2) can be problematic in the presence of two types of asymmetric response distributions. In one case, where most responses from the policy $\pi_{\theta}$ are high quality except for a single outlier, the relatively worse response may receive a negative advantage and contribute an opposing gradient, increasing variance. Although this response may still be correct under multi-dimensional reward rules, such as both format and correctness, it is penalized due to its format. This misalignment between reward attribution and the true value of a response increases the risk of reward hacking(Everitt et al., 2021; Pan et al., 2022). In the opposite case, where most responses are low quality and one is relatively better, the better one receives an excessively high advantage despite possibly low absolute quality, leading to biased updates and suboptimal convergence. AAPO augments advantage estimation with advantage momentum, addresses both issues: it boosts underappreciated yet valuable responses and suppresses misleading ones that exhibit a high estimated advantage. As a result, AAPO reduces variance and the risk of reward hacking, and improves the optimization stability of the policy.
Analysis 3  When policy optimization with AAPO reaches a global optimum, which means $\forall i,j\in\{1,2,3,\cdots,G\}\land i\neq j,\text{ }r_{i}\approx r_{j}$, the objective in equation 6, which consists of the CE loss and the augmented advantage, exhibits a specific and predictable behavior as detailed in our derivation below (for clarity and notational convenience, we omit the expectation operator). The optimization exhibits slight oscillations near the optimum, avoiding large gradient updates.
$\mathcal{J}_{\mathrm{AAPO}}(\theta)$ $\approx{\frac{1}{\sum_{i=1}^{G}\left|o_{i}\right|}\sum_{i=1}^{G}\sum_{t=1}^{\left|o_{i}\right|}\left(-\log\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)\cdot\operatorname{clip}(r_{\theta_{i}}-r_{ref_{i}},\delta_{\mathrm{low}},\delta_{\mathrm{high}})\right)}$
6 Experiments
6.1 Experimental setup
Most recent RL algorithms (Shao et al., 2024; Yu et al., 2025; Hu, 2025; Liu et al., 2025) struggle to make LLMs perform better at solving mathematical problems, which requires the model to think in the CoT (Wei et al., 2022) format before deciding the final answer. We choose open-rs (Dang & Ngo, 2025) as our training dataset for DeepSeek-R1-Distill-Qwen-1.5B base model, because the data in this dataset cover various types and difficulty levels of mathematical problems that are highly representative. To further provide a fair and rigorous evaluation of the effectiveness of our proposed AAPO, Qwen2.5-Math-7B model is chosen as the base model and subsequently trained on more challenging simplelr_qwen_level3to5 dataset (Zeng et al., 2025). In addition to the different model sizes, we also validate the effectiveness of AAPO on Llama series models.
In our experiment setting, we set the clip parameters $\delta_{\mathrm{low}}$ and $\delta_{\mathrm{high}}$ to be -0.2 and 0.28, respectively. We train all base models under the AAPO following the training process depicted in Algorithm 1. All rule-based reward functions adopted in our experiments are simple and straightforward. More training details about rule-based reward functions and the system prompt are provided in Appendix C. To evaluate the extent to which our proposed AAPO algorithm can enhance the reasoning capabilities of the model, we select AIME24, MATH-500 (Hendrycks et al., 2021; Lightman et al., 2024), AMC23, Minerva (Lewkowycz et al., 2022), and OlympiadBench (Huang et al., 2024) as evaluation benchmarks. Our evaluation framework utilizes a well-established and community-vetted codebase, maintaining consistency with widely adopted implementations.
6.2 Results
As shown in Table 1, taking model DeepSeek-R1-Distill-Qwen-1.5B as our base model, the application of our proposed AAPO enables the base model to achieve the SOTA performance on Minerva, the second-best performance on MATH-500, AMC23 and OlympiadBench. When averaging scores across all benchmarks, the resulting AAPO-1.5B model achieves an overall SOTA performance. By direct comparison under the same training data, AAPO-1.5B achieves improvements of 2.7% and 2.0% over GRPO-1.5B and GPG-1.5B, respectively. It is worth noting that our AAPO-1.5B achieves performance superior to Still-3-1.5B-Preview (Team, 2025), despite the fact that Still-3-1.5B-Preview benefits from a larger training dataset that consists of long CoT reasoning data distilled from the DeepSeek-R1 (Guo et al., 2025) model and performs RL training with specified reward strategies. After employing our proposed AAPO on the Qwen2.5-Math-7B model, AAPO-7B achieves the overall SOTA performance compared to other methods. AAPO-7B achieves improvements of 12.2%, 10.3% over SimpleRL-Zero-7B (Zeng et al., 2025), GPG-7B (Chu et al., 2025), respectively, under identical training data. Furthermore, AAPO-7B also outperforms other models such as Eurus-2-7B-PRIME (Cui et al., 2025), OpenReasoner-Zero-7B (Hu et al., 2025), despite the fact that these baselines are trained with more high-quality data or data distilled from the DeepSeek-R1 model. This suggests that the effectiveness can be mainly attributed to the design of our proposed AAPO rather than to the scale or quality of the training data. AAPO also demonstrates superior performance compared to GRPO and GPG on Llama series models under the same training and evaluation settings as reported in Table 1. Compared to GRPO and GPG, AAPO achieves absolute improvements of 3.5%, 2.5%, and 4.6%, 2.6% on Llama 1B and 3B models, respectively. It is worth noting that evaluation results may be influenced by the computational device type. We have also provided the original results of each model from the corresponding papers in Appendix E.
Model AIME24 MATH-500 AMC23 Minerva OlympiadBench Avg.
Llama 1B Models
Llama-3.2-1B-Instruct 0.0 11.8 2.5 1.8 3.7 4.0
+GRPO (Shao et al., 2024) 0.0 19.4 12.5 3.7 4.3 8.0
+GPG (Chu et al., 2025) 0.0 21.2 17.5 1.8 4.7 9.0
[…]
Llama 3B Models
Llama-3.2-3B-Instruct 3.3 28.6 7.5 3.7 7.6 10.1
+GRPO (Shao et al., 2024) 0.0 32.8 22.5 8.1 7.7 14.2
+GPG (Chu et al., 2025) 6.7 40.0 15.0 8.8 11.6 16.2
[…]
Qwen 1.5B Models
DeepSeek-R1-Distill-Qwen-1.5B 33.3 84.4 70.0 30.9 50.8 53.9
GRPO-1.5B (Dang & Ngo, 2025) 26.7 86.2 82.5 27.6 52.6 55.2
GPG-1.5B (Chu et al., 2025) 36.7 83.4 75.0 29.8 53.2 55.6
Still-3-1.5B-Preview (Chen et al., 2025) 40.0 85.5 72.5 30.5 53.9 56.5
[…]
Qwen 7B Models
Qwen2.5-Math-7B (Yang et al., 2024) 6.7 56.2 47.5 14.0 23.4 39.6
SimpleRL-Zero-7B (Zeng et al., 2025) 30.0 77.4 57.5 30.5 38.1 46.7
GPG-7B (Chu et al., 2025) 23.3 80.2 55.0 36.0 42.8 47.5
OpenReasoner-Zero-7B (Hu et al., 2025) 20.0 80.8 65.0 29.4 46.2 48.3
Eurus-2-7B-PRIME (Cui et al., 2025) 16.7 81.8 65.0 37.5 44.6 49.1
Oat-Zero-7B (Liu et al., 2025) 30.0 81.2 65.0 34.9 43.4 50.3
[…]
Table 1: Zero-shot pass@1 performance on mathematical reasoning benchmarks. represents reproduced results with our best effort under the same settings. Bold and Underline indicate the best and the second-best performance in the corresponding category, respectively.
6.3 Ablation study
Clip operation To investigate the contribution of the clip operation to AAPO, we conduct additional ablation studies by removing the clip operation on both the 1.5B and 7B models. The results presented in Table 2 indicate that the performance of the optimized model without clip is inferior to that with the clip, suggesting that incorporating the clip operation effectively contributes to further improvements in optimization results. As illustrated in the right figure, the optimization process becomes more stable with the incorporation of the clip operation compared to the optimization process without it. The resulting AAPO-7B exhibits better performance on the benchmarks when the clip operation is adopted.
Model AIME24 MATH-500 AMC23 Minerva OlympiadBench Avg.
AAPO-1.5B 33.3 86.0 80.0 30.9 53.3 56.7
AAPO-1.5B w/o clip 33.3 85.0 82.5 29.0 53.3 56.6
AAPO-7B 30.0 82.4 70.0 35.3 44.3 52.4
AAPO-7B w/o clip 30.0 79.6 70.0 34.9 42.5 51.2
Training process analysis  As illustrated by the training curves in Figure 2, AAPO achieves optimization performance comparable to that of GPG, while exhibiting further improvements during the later steps of RL training. In the Format Reward and Reward figures, AAPO consistently attains higher reward than GPG in the later steps of the training. In addition, three Reward figures demonstrate significantly reduced fluctuations during the training process. AAPO also outperforms GRPO. Moreover, in the Response Length figure, our proposed AAPO demonstrates superior optimization, indicating higher training efficiency. The final results presented in Table 1 demonstrate that our proposed AAPO achieves superior performance across five mathematical reasoning benchmarks. To straightly and effectively analyze the training stability of AAPO, we calculated the variance of the training loss forAAPO as well as methods GRPO and GPG. $\operatorname{Var}(\mathcal{L}_{\mathrm{AAPO}})=3.6\times 10^{-4}$, $\operatorname{Var}(\mathcal{L}_{\mathrm{GPG}})=3.9\times 10^{-4}$ and $\operatorname{Var}(\mathcal{L}_{\mathrm{GRPO}})=3.53\times 10^{-3}$. The results show that AAPO exhibits the lowest variance, indicating relatively stable training. Additionally, the results of the ablation experiments on $\operatorname{clip}$ operation demonstrate that AAPO can steadily improve model performance as training epochs progress. More experiment analysis can be found in Appendix D.
Figure 2: Training process of DeepSeek-R1-Distill-Qwen-1.5B on open-rs (Dang & Ngo, 2025) utilizing our proposed AAPO algorithm. Compared to GPG (Chu et al., 2025) and GRPO (Shao et al., 2024), AAPO demonstrates better stability during training and achieves superior performance in the final results as shown in Table 2.
7 Conclusion
In this paper, we conduct an in-depth analysis of the limitations inherent in the group relative advantage estimation method used by mainstream RL algorithms, such as GRPO, which would lead to optimization issues such as zero gradient and gradient ascent. To address these issues, we propose a novel RL algorithm Advantage-Augmented Policy Optimization (AAPO). By augmenting the group relative advantage estimation method with advantage momentum, our method effectively improves policy optimization performance in experimental benchmarks. Experimental results across several mathematical reasoning benchmarks and model series demonstrate that AAPO achieves the overall superior performance across several mathematical reasoning benchmarks.
Appendix A Algorithm
@@
AAPO Training The training process of AAPO can be summarized as follows: 1) Sample two groups of responses, $O_{\theta}$ and $O_{ref}$, of equal size from both the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$, respectively. 2) For each response in $O_{\theta}$, compute its group relative advantage and the advantage momentum with the corresponding response in $O_{ref}$. The values, calculated by equation 5, are then used to perform gradient updates according to equation 6. It is important to note that during the training process, the parameters of the reference model $\pi_{ref}$ remain frozen and do not undergo gradient updates. The training procedure is described in Algorithm 1.
@@
8:   5) compute the augmented advantage $\hat{A}_{i,t}^{*}$ following equation 5 for each sample in $O_{\theta}$
9:   6) compute loss for $O_{\theta}$ following equation 4
@@
11:  7) update $\pi_{\theta}$ following equation 6
@@
Appendix B Proof
Theorem 1. (Stability)  Since the rewards are bounded, the group standard deviation satisfies $0\leq\sigma_{min}\leq\sigma$, and the log-likelihood gradients are bounded as $\nabla_{\theta}\log\pi_{\theta}(o)||\leq M$. Then, Each gradient step with learning rate $\eta_{k}$ satisfies $\theta_{k+1}-\theta_{k}||\leq\eta_{k}MB$, where $B=\frac{R_{max}-R_{min}}{\sigma_{min}}+\operatorname{max}(|\delta_{\mathrm{low}}|,|\delta_{\mathrm{high}}|)$ is the uniform bound on the AAPO weights. The expected objective is bounded from $\mathcal{L}(\theta)\geq-B\log|\mathcal{V}|$, where $|\mathcal{V}|$ is the vocabulary size. Hence, AAPO training is stable: the objective cannot diverge to -$\infty$ and parameter updates are always finite.
Proof.
We restate the definitions used: (1) For any response $o$, the advantage $\hat{A}^{*}$ satisfies $|\hat{A}^{*}|\leq B=\frac{R_{max}-R_{min}}{\sigma_{min}}+\operatorname{max}(|\delta_{\mathrm{low}}|,|\delta_{\mathrm{high}}|)$, where $\sigma_{min}>0$ is the minimal standard deviation ensured by group sampling. (2) The gradient of the log-policy is bounded: $\nabla_{\theta}\log\pi_{\theta}(o)||\leq M$, and $\log\pi_{\theta}(o)$ is $L_{0}$-Lipschitz. (3) The vocabulary size satisfies $|\mathcal{V}|<\infty$. For a group $\mathcal{G}$, the gradient of the empirical loss is:
$\nabla_{\theta}\mathcal{L}_{\mathcal{G}}(\theta)=\frac{1}{N_{\mathcal{G}}}\sum_{o\in O}\nabla_{\theta}[-\log\pi_{\theta}(o)]\hat{A}^{*}.$
By assumption (1) and (2), we have
$\nabla_{\theta}\mathcal{L}_{\mathcal{G}}(\theta)||\leq\frac{1}{N_{\mathcal{G}}}\sum_{o\in O}M\cdot B=MB.$
Thus, one gradient update with learning rate $\eta_{k}$ yields
[…]
For any response $o$,
$\mathbb{E}_{a\sim\pi_{\theta}}\left[-\log\pi_{\theta}(o)\right]=H(\pi_{\theta}(\cdot\mid o_{<t}))\leq\log|\mathcal{V}|.$
where $H(\cdot)=-\sum_{x}p(x)\log p(x)$ denotes Shannon entropy (Shannon, 1948). Multiplying by the bounded advantage $|\hat{A}^{*}|\leq B$ and averaging over tokens in the batch, we obtain
$\mathcal{L}(\theta)\geq-B\log|\mathcal{V}|.$
we conclude that parameter updates are bounded and the objective is lower bounded (which means the objective cannot diverge to -$\infty$). Therefore, AAPO training is stable and cannot diverge. ∎
Theorem 2. (Convergence)  Assume that the stochastic gradient is unbiased and that the per-sample gradient has bounded second moment. Let the step sizes satisfy the Robbins–Monro conditions $\eta_{k}>0$, $\sum_{k}\eta_{k}=\infty$, $\sum_{k}\eta_{k}^{2}<\infty$. AAPO converges to a stationary point of its expected objective $\liminf_{k\to\infty}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]=0$. Moreover, if a constant step size $\eta<\frac{1}{BL_{0}}$ is used, where $L_{0}$ is the smoothness constant of $\log\pi_{\theta}(o)$, then the iterates converge to a neighborhood of stationarity $\limsup_{K\to\infty}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]\;\lesssim\;\mathcal{O}(\eta)+\mathcal{O}\!\left(\tfrac{1}{N_{\mathcal{G}}}\right)$.
Proof.
We adopt the same definitions as in the Proof of Theorem 1. Since $\log\pi_{\theta}(o\mid q)$ is $L_{0}$-smooth (i.e., its gradient is $L_{0}$-Lipschitz) and the advantage is bounded by $B$, the empirical loss $\mathcal{L}_{\mathcal{G}}(\theta)$ is $L=BL_{0}$-smooth:
$\left\|\nabla\mathcal{L}_{\mathcal{G}}(\theta)-\nabla\mathcal{L}_{\mathcal{G}}(\theta^{\prime})\right\|_{2}\leq L\left\|\theta-\theta^{\prime}\right\|_{2}.$
By the descent lemma for $L$-smooth functions, for any step size $\eta_{k}$,
$\mathcal{L}_{\mathcal{G}}(\theta-\eta_{k}g)\leq\mathcal{L}_{\mathcal{G}}(\theta)-\eta_{k}\langle g,\nabla\mathcal{L}_{\mathcal{G}}(\theta)\rangle+\tfrac{L}{2}\eta_{k}^{2}\|g\|^{2}.$
Choosing $g=\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})$ and requiring $\eta_{k}\leq 1/L$, we obtain
$\mathcal{L}_{\mathcal{G}}(\theta_{k+1})\leq\mathcal{L}_{\mathcal{G}}(\theta_{k})-\tfrac{\eta_{k}}{2}\left\|\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})\right\|^{2}.$
Assume that, for all $\theta$, the per-sample loss $\ell(q,o;\theta)=-\log\pi_{\theta}(o\mid q)\,\hat{A}^{*}$ has a gradient with bounded second moment:
$\sup_{\theta}\mathbb{E}_{(q,o)\sim\pi_{\theta}}\!\left[\left\|\nabla_{\theta}\ell(q,o;\theta)\right\|^{2}\right]\leq\sigma^{2},$
where $\sigma^{2}$ is a constant. Since $\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})$ is the average of $N_{\mathcal{G}}$ i.i.d. samples conditional on $\theta_{k}$, we have